ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 81, 1993

Acosta, M. D. Every Banach space can be renormed to satisfy the denseness of numerical radius attaining operators

273

Argyros, S. A. and I. Deliyanni Non-dentable sets in Banach spaces with separable dual

53

Bayer, M. M. Equidecomposable and weakly neighborly polytopes  301

Bezdek, K., Gy. Kiss and M. Mollard An illumination problem for zonoids

265

Camps, R. and W. Dicks On semilocal rings  203

Castillo, J. M. F. and M. González The Dunford–Pettis property is not a three-space property

297

Chen, J.-C. Weights and L_{Φ}-boundedness of the Poisson integral operator

193

Chiappinelli, R. and D. E. Edmunds Eigenvalue asymptotics and a non-linear Schrödinger equation

179

deLaubenfels, R. C-semigroups and strongly continuous semigroups

227

Deliyanni, I. See S. A. Argyros

Dicks, W. See R. Camps

Edmunds, D. E. See R. Chiappinelli

Feldman, J. A generalization of a result of R. Lyons about measures on [0,1)

281

Glasner, E. Minimal nil-transformations of class two  31

González, M. See J. M. F. Castillo

Harpe, de la P., A. G. Robertson and A. Valette On the spectrum of the sum of generators for a finitely generated group

65

Herrero, M. A. and J. J. L. Velázquez Plane structures in thermal runaway

321

Juhász, I. On the weight-spectrum of a compact space  369
Kella, O. Star search—a different show  145

Kemer, A. R. The standard identity in characteristic p: a conjecture of I. B. Volichenko 343 Kirkman, E. E. and L. W. Small q-Analogs of harmonic oscillators

and related rings

111

Kiss, Gy. See K. Bezdek

Landver, A. Finite combinations of Baire numbers  289

Liskevich, V. A. and E. W. Tuv A priori estimates for second order elliptic equations and their applications

257

Loveys, J. and Y. Peterzil Linear o-minimal structures  1
Mekler, A. H. and S. Shelah Every coseparable group may be free 161

Mollard, M. See K. Bezdek

Nagel, R. and F. Räbiger Superstable operators on Banach spaces 213
Nielsen, M. J. On two questions concerning tilings  129
Peck, N. T. Twisted sums and a problem of Klee  357

Peterzil, Y. See J. Loveys Räbiger, F. See R. Nagel Robertson, A. G. See P. de la Harpe Shelah, S. and L. Stanley More consistency results in partition

calculus

97

Shelah, S. See A. H. Mekler Small, L. W. See E. E. Kirkman Stanley, L. See S. Shelah Tuv, E. W. See V. A. Liskevich Valette, A. See P. de la Harpe Velázquez, J. J. L. See M. A. Herrero